OpenStudy (anonymous):

what is the value of 3m+2n if m=5 and n=3?

OpenStudy (anonymous):

3(5) + 2(3) = 15 +6 = 21

OpenStudy (anonymous):

find the width of a rectangle with the length 6.4mm abd area 20.48mm